OpenStudy (anonymous):

Verify the identity. 4 csc 2x = 2 csc^2x tan x

OpenStudy (anonymous):

sin ( 2x) = 2 sin (x) cos (x)

OpenStudy (anonymous):

csc (x) = 1/ sin (x)

OpenStudy (anonymous):

Where did the 4 go?

OpenStudy (anonymous):

it turned into a 2

OpenStudy (anonymous):

\[4 \csc(2x) = \frac{4}{\sin(2x)}\]

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

do you know how to solve it now or do i need to walk you through it?

OpenStudy (anonymous):

I understand now. It's a double angle identity

OpenStudy (anonymous):

ok